OpenStudy (anonymous):

Simplify the expression (4x^5)(5x^6).

OpenStudy (bibby):

\(\huge a^m*a^n=a^{m+n}\)

OpenStudy (anonymous):

Oh i just add the exponents?

OpenStudy (bibby):

for the x's. remember that multiplication is commutative so a*b*c=c*a*b as such we can rewrite the expression \((4x^5)(5x^6)=(4*5)(x^5*x^6)\)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

can you show me step by step how you got that please? >.<

OpenStudy (bibby):

hmm, let's see. all I did was rearrange the terms. step 1: \((4x^5)(5x^6)=4*x^5*5*x^6\) you can rearrange all the terms because of the commutativity of multiplication step 2: \(4*x^5*5*x^6=4*5*x^5*x^6\) step 3: \((4*5)(x^5*x^6)\) step 4: simplifying the above

OpenStudy (bibby):

so far I haven't really done anything

OpenStudy (bibby):

just rearranged numbers/terms

OpenStudy (bibby):

do you need another push?

OpenStudy (anonymous):

Yea can you simplify it?

OpenStudy (bibby):

what is 4*5? what is x^5*x^6?

OpenStudy (anonymous):

4*5=20

OpenStudy (bibby):

yeah

OpenStudy (bibby):

what about the other thing

OpenStudy (anonymous):

multiply the exponents?

OpenStudy (anonymous):

my math teacher is gonna kill me T^T

OpenStudy (bibby):

that's illegal your math teacher is probably an alright guy or gal just look at my first post to recall the relevant property of exponents

OpenStudy (bibby):

\({a^b}^c=a^{b*c}\)

OpenStudy (bibby):

but here we're adding the exponents

OpenStudy (bibby):

because it's multiplication of powers

OpenStudy (bibby):

I'm pretty sure I got my words backwards. whatever, when you multiply things with exponents and a common base, you add the powers

OpenStudy (anonymous):

ok

OpenStudy (bibby):

a^m*a^n=a^(m+n)

OpenStudy (bibby):

\(a^m*a^n=a^{(m+n)}\)

OpenStudy (anonymous):

lol...

OpenStudy (bibby):

what's so funny